The stopcock connecting a 1.12 L bulb containing helium gas at a pressure of 6.28 atm, and a 2.32 L bulb containing carbon dioxide gas at a pressure of 3.11 atm, is opened and the gases are allowed to mix. Assuming that the temperature remains constant, the final pressure in the system is  atm.
Answer #1

SOLUTION :


P*V of helium =  = 6.28 * 1.12 = 7.0336 L-atm

P*V of CO2 = 3.11 * 2.32 = 7.2152  L-atm


When these two bulbs are connected through a cock  and allowed to mix and temperature remains constant.


P*V of the system = Total of the P*V of the two bulbs = 7.0336 + 7.2152 = 14.2488 L-atm


=> P of the system * V of the system = 14.2488

=> P of the system = 14.2488 / V of the system

=> P of the system = 14.2488 / (1.12 + 2.32)

=> P of the system = 4.1421  atm 


Final pressure in the system is 4.1421 atm (ANSWER)

answered by: Tulsiram Garg